Harris wants to sell his share of an investment to Bob for $50,000 in two years. If money is worth 12% compounded quarterly, what would Harris accept today?

a. $39,471.

b. $41,874.

c. $11,941.

d. $41,000.
Answer #1

Answer : S 39471

Annual rate = 12 %

Quarterly rate = 12 /4 = 3 %

For two years = 8 Quarters

Price to accept today = [ $ 50000 / ( 1 + 0.03 )8]

Price to accept today = [ $ 50000 / 1.2668 ] = $ 39471